OpenStudy (anonymous):

converge or diverge {sum_{1}^{infty} 5^{-3} pi ^{k}}

OpenStudy (experimentx):

Looks much better. \[ {\sum_{k=1}^{\infty} 5^{-3} \pi ^{k}}\] Clearly it diverges since it series from geometric progression of common ratio \(\pi\)

OpenStudy (anonymous):

okay thanks. just making sure

OpenStudy (experimentx):

you are welcome.